OpenStudy (nuccioreggie):

Conner and Jana are multiplying (3568)(39610). Conner's Work Jana's Work (3568)(39610) = 35 + 968 + 10 = 314618 (3568)(39610) = 35⋅968⋅10 = 345680 Is either of them correct? Explain your reasoning.

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

Please help

OpenStudy (anonymous):

I'm gonna assume this is using exponents. Am I right?

OpenStudy (nuccioreggie):

Yes and here is what i got They are both incorrect. There results should be way more then in the thousands range, they should be more like in the millions.

OpenStudy (anonymous):

Do you mind drawing out what the exponents look like so I can see?

OpenStudy (nuccioreggie):

Here it is

Nnesha (nnesha):

please next time type your question don't just copy&paste it's also against the COC or TAC :-) :-)

OpenStudy (nuccioreggie):

Ok but is my answer right

OpenStudy (nuccioreggie):

??

Nnesha (nnesha):

mhm i was waiting 4 legend reply well what's the exponent rule ? when you multiply same bases what you should do with their exponents ?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @nuccioreggie Yes and here is what i got They are both incorrect. There results should be way more then in the thousands range, they should be more like in the millions. \(\color{blue}{\text{End of Quote}}\) nope. you don't need a calculator solve this question

Nnesha (nnesha):

to*

OpenStudy (michele_laino):

please use this rule: \[\Large {a^m} \times {a^n} = {a^{m + n}}\]

TheSmartOne (thesmartone):

Your answer is wrong. Use the formula Michele gave and work from there.

Nnesha (nnesha):

write these dwn 'n ur notebook exponent rules \[\large\rm x^m \times x^n = x^{m +n}\]\[\large\rm (x^m)^n = x^{m \times n}\]\[\large\rm \frac{ x^m }{ x^n }=x^{m-n}\]